# Sentence Smith: Formally Controllable Text Transformation and its Application to Evaluation of Text Embedding Models

### v2

@@ Abstract
We propose the SentenceSmith framework that enables controlled and specified manipulation of text meaning. It consists of three main steps: 1. Parsing a sentence into a semantic graph, 2. Applying human-designed semantic manipulation rules, and 3. Generating text from the manipulated graph. A final filtering step (4.) ensures the validity of the applied transformation. To demonstrate the utility of SentenceSmith in an application study, we use it to generate hard negative pairs that challenge text embedding models. Since the controllable generation makes it possible to clearly isolate different types of semantic shifts, we can gain deeper insights into the specific strengths and weaknesses of widely used text embedding models, also addressing an issue in current benchmarking where linguistic phenomena remain opaque. Human validation confirms that the generations produced by SentenceSmith are highly accurate.
Sentence Smith: Formally Controllable Text Transformation and its Application to Evaluation of Text Embedding Models
Hongji Li Andrianos Michail Reto Gubelmann Simon Clematide Juri Opitz University of Zurich {hongji.li,andrianos.michail,reto.gubelmann,simon.clematide}@uzh.ch opitz.sci@gmail.com
@@ 1 Introduction
Transforming a sentence’s meaning through computational methods is a challenging task with numerous applications, including style transfer, data augmentation, and controlled text generation. Large language models (LLMs) offer a promising approach via prompting but lack fine-grained control (Greenblatt et al., 2024; Mizrahi et al., 2024). Conversely, symbolic methods, e.g., word replacement using taxonomy lookups (Bolshakov and Gelbukh, 2004; Huang et al., 2009), provide control but often produce unnatural sentences with limited variation.
@@
To bridge the gap between these paradigms, we propose SentenceSmith, a neuro-symbolic framework for text manipulation. The process begins with a parser that maps a sentence onto a symbolic graph-based meaning representation, such as an Abstract Meaning Representation (AMR) (Banarescu et al., 2013). We posit that such a representation provides an effective interface for applying well-defined, precise, and targeted meaning-altering operations. Once the graph is updated, a generator converts it back into natural text.
@@
This process is illustrated in Figure 1, where a single sentence is transformed into two new sentences with distinct meanings. On the left, we reverse semantic roles (e.g., tiger, snake), modifying the event structure of the sentence by swapping patient and agent roles. On the right, we apply a two-step process: removing the patient (tiger) and negating the main predicate (bite-01), resulting in a new sentence that describes a “harmless” snake (the “-01” in bite-01 indicates a specific sense in PropBank (Palmer et al., 2005)). We also introduce an optional faithfulness check as a post-processing step to evaluate the consistency between the transformed graph and the generated sentence, filtering out outputs affected by parsing or generation errors.
@@
We argue that this combination of symbolic control, neural fluency, and post-hoc verification provides a powerful and flexible approach to sentence meaning transformation, offering a more controllable alternative to prompting LLMs.
@@
To illustrate the practical value of SentenceSmith, we conduct a detailed demonstration study, leveraging our framework to generate hard negative pairs that pose significant challenges for text embedding models. By applying well-defined transformation rules targeting specific semantic phenomena, SentenceSmith creates nuanced sentence pairs that go beyond superficial lexical differences, enabling a more fine-grained evaluation of embedding models. Through this application study, we shed light on their strengths and weaknesses in handling particular linguistic phenomena, addressing a limitation in current benchmarking practices, where such intricacies are often obscured.
@@
Broadly speaking, this study contributes to the ongoing effort to pinpoint the limits of Transformer-Based Neural Networks’ understanding of sentence meaning. It has been established that even the latest State-of-the-Art LLMs struggle with important sentence-level meaning phenomena such as negation (Parmar et al., 2024), and Xu et al. (2025) finds that LLMs still struggle with exact, syntax-based modes of reasoning, sometimes being outperformed by BART (Lewis et al., 2020), an early sequence-to-sequence transformer. The fine-grained and dynamic experimental settings enabled by SentenceSmith shed new light on this central, but still open question regarding the abilities of Transformers.
@@
The remainder of this paper is structured as follows: Section §2 reviews background and related work. Section §3 introduces our neuro-symbolic SentenceSmith framework, detailing its components, including the parser, symbolic graph transformations, generator, and optional faithfulness checker. Section §4 presents an application demonstration, focusing on generating hard negative pairs to challenge state-of-the-art text embedding models and reveal potential weaknesses in their linguistic understanding. Finally, Section §5 concludes the paper, discussing the broader implications of our work and potential directions for future research. We release code and data under a public license.111https://github.com/1160300921/SentenceSmith
@@ 2 Related Work and Background
The value of using semantic graphs as an intermediate representation, particularly AMR, has been highlighted in two recent surveys (Sadeddine et al., 2024; Wein and Opitz, 2024). This approach allows the fusion of neural network power with the expressivity and explicitness of meaning representations, which is especially useful when interpretability and control are required in an application. Related work applies this principle to style transfer (Jangra et al., 2022), translationese reduction222Translationese reduction aims at attempting to reduce “language weirdness” that can arise when translating texts, the authors show that post-processing with parsing and generation can mitigate these artifacts. (Wein and Schneider, 2024), and data augmentation (Shou et al., 2022; Shou and Lin, 2023; Ghosh et al., 2024). Our work generalizes this principle further and imposes an additional check for validating the faithfulness of the generation. Notably, the idea of planning and controlling sentence generation through meaning representation dates decades back (i.a., Sondheimer and Nebel, 1986; Mann and Nigel, 1986; Kasper, 1989; Wijnen, 1990; Bateman et al., 1990), but was limited by inaccuracies parsing and generation systems. With stronger parsing and generation systems now at hand, we showcase the usefulness of this way of controllable text generation.
@@
Text embedding models are crucial for a wide range of NLP tasks, including semantic search, information retrieval, and NLG evaluation (Clark et al., 2019; Muennighoff, 2022; Gao et al., 2023). Since Reimers and Gurevych (2019)’s foundational "SBERT" work, multiple branches of embedding model research have emerged. These include enhancing model performance through scaling parameters (Wang et al., 2023) or training data (Wang et al., 2022a), as well as exploring unsupervised embeddings (Gao et al., 2021) and interpretable embeddings (Opitz and Frank, 2022b). However, key questions remain: What is the accuracy of such embeddings? What level of linguistic understanding resides in those vectors? Towards being better able to answer such questions, we employ SentenceSmith and demonstrate its capacity to generate fine evaluation data that test embedding models’ ability to assess different linguistic phenomena. In posing such questions towards better understanding model’s similariy decisions through assessing behavior on tailored datasets, our goal is related to a recent/concurrent line of work from Nastase et al. (Nastase and Merlo, 2024; Nastase et al., 2024b, a).
Then, we would also like to learn more about different notions of similarity: Even for long-established datasets like SICK (Marelli et al., 2014) or STS (Agirre et al., 2016), it remains unclear what specific aspects human annotations elicit. For instance, Budanitsky and Hirst (2006) point out that while relatedness (SICK) and similarity (STS) are somewhat related notions, similarity is “not an adequate proxy (for relatedness).” Recent empirical research investigates sentence meaning through descriptions, assessing the similarity of texts through these descriptions (Hoyle et al., 2023; Ravfogel et al., 2024). Likewise, paraphrases vary widely and are often difficult to formally distinguish (Michail et al., 2025). In our work, we formalize relatedness as a confounder of similarity by inducing pairs with varying degree of structural surface similarity, while their potential meaning differences are formally distinguishable. Furthermore, the stasis of most benchmarks has also drawn upon criticism (Fan et al., 2024), and limits evaluation to the available data, with potential ramifications for the trustworthiness of results. By demonstrating how SentenceSmith can generate challenging, trustworthy, and interpretable test sets, we pave the way for more customizable, dynamic and interpretable testing of models.
@@ 3.1 Model Overview
where are graph transformations that we apply on the graph representation of the input sentence , output of . Finally, is used to generate a text from the manipulated graph.
@@ 3.2 Parameterization
We rely on Abstract Meaning Representation (AMR, Banarescu et al., 2013) as our graph framework. AMR represents text as directed acyclic graphs, where nodes denote entities and edges capture semantic relations. The overarching goal of AMR is to explicitly encode “Who does what to whom?”. This explicitness is a key motivation for our work, as it enables targeted meaning changes within the semantic structure.
@@
As is standard in most AMR applications, we employ off-the-shelf parsing and generation models.333https://github.com/bjascob/amrlib. These models, based on pre-trained BART (Lewis, 2019), produce linearized sequence graphs from text (parsing) or text from linearized sequence graphs (generation). To facilitate manipulation, we convert the linearized graph into triples of the form <u,r,v>, where u and v are graph nodes, and r is a relation.
@@
The criteria for discarding output sentences depend on the application: If the goal is meaning alteration, we should filter out cases where and mutually entail each other. Conversely, if we seek paraphrases, we should discard cases where contradiction or neutrality is detected. SentenceSmith parameterizes with an efficient NLI-based model from Steen et al. (2023). This system is robustness-enhanced through data augmentation and achieves strong results on the TRUE faithfulness benchmark (Honovich et al., 2022).444We use https://huggingface.co/juliussteen/DeBERTa-v3-FaithAug.
@@ 4 Application Study
The foil concept is illustrated in Figure 2. The objective is to generate a new sentence that is structurally highly similar to an existing one but semantically distinct. If this transformation is applied to one part of a paraphrase pair, the resulting sentence will not maintain a paraphrase relation with either of the original sentences. In the example, we start with the paraphrases a) A tiger is being bitten by a serpent. and its paraphrastic variant b) The snake bites the tiger. Our parser processes sentence b), and within the generated meaning space, we perform a semantic role confusion, swapping the agent and patient of the bite event. Finally, our generator produces c) The tiger bites the snake, which is structurally more similar to a) than a) and b) are to each other (i.e., higher lexical overlap). However, the event’s semantics have fundamentally changed, thereby breaking the paraphrase relation.
@@ 4.1.1 Defining Semantic Manipulations
This fundamental manipulation assesses embedding models’ sensitivity to negation. We define negation as altering the polarity of a predicate within an AMR graph. Specifically, this transformation attaches a <n,:polarity,-> edge (in AMR notation, a negation) to a randomly selected node n, excluding pronouns. For example, modifying the predicate approve by adding this edge results in its opposite, disapprove (i.e., in AMR: <n,:polarity,-> is added to node n, where <n,:instance,approve> already exists). The manipulated AMR is then used to generate the negated text. Also in the example of Figure 3, the truth value of the sentence changes.
2. Role Swap (RS)
3. Underspecification (US).
@@ 2. Role Swap (RS)
This transformation removes a randomly selected leaf node n from the AMR graph, reducing semantic content, thereby breaking a paraphrase relation. The introduced ambiguity or incompleteness tests how robust embedding models are to partial information.555Note that it can also introduce a polarity change, in case a negation node in the AMR is removed. The notion of “Underspecification” refers to the AMR structure being reduced. In Figure 3, the patient specification (a snake) has disappeared.
4. Antonym Replacement (AR)
@@
inverts the meaning of a selected node while maintaining the context of the surrounding graph. The system identifies a random non-pronoun node, extracts its lexical stem (e.g., happy in happy-01), and queries WordNet (Miller, 1995) for antonyms. If a suitable antonym is found, the node label is updated while preserving any sense-suffix (e.g., -01). This transformation alters meaning more explicitly than polarity negation by directly substituting an opposite term. In Figure 3, same as in PN, the truth value changes —while still a biting is happening, this time the agent is a small rather than big.
5. Hypernym Substition (HS).
@@
This transformation replaces a word with its hypernym, distorting inferences and breaking paraphrase relationships. Consider “Penguins can’t fly”. If Penguin is replaced with its hypernym “Bird”, the sentence becomes nonsensical (“Birds can’t fly”), demonstrating how hypernym substitution alters meaning while preserving surface structure. Following the same WordNet-based approach as in the antonym replacement, non-pronoun node labels are substituted with randomly selected hypernyms. The node’s graph position and semantic roles remain unchanged, but abstraction increases, disrupting logical inferences. In Figure 3, we observe how snake is replaced with animal. This is also some form of underspecification, but speaking through the AMR, it is achieved by replacing (not removing) a concept (which can also trigger a changed truth value, as in the aforementioned bird-example).
4.1.2 Discussion
While introducing more control in changing text meaning, it is likely that our manipulations not always yield a desired outcome. Clearly, we want to prevent the generation of a paraphrase, or near-paraphrase. We do not expect that this happens often, since all manipulations are bound to change the meaning in some way or another; and even parsing or generation errors are much more likely to generate something that is not a paraphrase of the input than otherwise. But generally, there may be several less desirable outcomes. One of these would be the generation of weird sentence meanings (e.g., swapping the roles of horse and leaves in the horse eats leaves would yield such a weird meaning). Another less desirable outcome may be very similar sentence meanings (e.g., if a hypernym is semantically very close to its hyponym). While an argument can be made for the former that these weird sentences are still fully valid foils for embedding models (weird meaning or not, the paraphrase relation is broken), the latter may be the least desirable type (since they may be close to actual paraphrases). This is where our entailment check proves valuable, since it allows us to fish out those foils that we can trust in the most (e.g., the contradictory ones).

### v4

# Sentence Smith: Controllable Edits for Evaluating Text Embeddings

@@ Abstract
Controllable and transparent text generation has been a long-standing goal in NLP. Almost as long-standing is a general idea for addressing this challenge: Parsing text to a symbolic representation, and generating from it. However, earlier approaches were hindered by parsing and generation insufficiencies. Using modern parsers and a safety supervision mechanism, we show how close current methods come to this goal. Concretely, we propose the SentenceSmith framework for English, which has three steps: 1. Parsing a sentence into a semantic graph. 2. Applying human-designed semantic manipulation rules. 3. Generating text from the manipulated graph. A final entailment check (4.) verifies the validity of the applied transformation. To demonstrate our framework’s utility, we use it to induce hard negative text pairs that challenge text embedding models. Since the controllable generation makes it possible to clearly isolate different types of semantic shifts, we can evaluate text embedding models in a fine-grained way, also addressing an issue in current benchmarking where linguistic phenomena remain opaque. Human validation confirms that our transparent generation process produces texts of good quality. Notably, our way of generation is very resource-efficient, since it relies only on smaller neural networks.
Sentence Smith: Controllable Edits for Evaluating Text Embeddings
Hongji Li Andrianos Michail Reto Gubelmann Simon Clematide Juri Opitz University of Zurich Correspondence: opitz.sci@gmail.com
@@ 1 Introduction
How can we transform the meaning of a sentence such that the output remains fluent while the transformation process is maximally transparent? Some symbolic methods, e.g., those based on word replacement using taxonomy lookups (bolshakov2004synonymous; huang2009using), provide transparency and control but often produce unnatural sentences with limited variation. On the other end, of course, we now have the all-dominating paradigm of LLM prompting. But clearly this process is opaque, and any control is, at best, indirect (greenblatt2024ai; 10.1162/tacl_a_00681).
@@
To bridge the gap between these paradigms, and to highlight an alternative to LLM-based methods, we propose SentenceSmith, a neuro-symbolic framework for text manipulation. The process begins with a parser that maps a sentence onto a symbolic graph-based meaning representation, such as an Abstract Meaning Representation (AMR, banarescu2013abstract). We posit that such a representation provides an effective interface for applying well-defined, precise, and targeted meaning-altering operations. Once the graph is updated, a generator converts it back into natural text.
@@
The process is illustrated in Figure 1, where a single sentence is transformed into two new sentences with different meanings. On the left, we reverse semantic roles (e.g., tiger, snake), modifying the event structure of the sentence by swapping patient and agent roles. On the right, we apply a two-step process: removing the patient (tiger) and negating the main predicate (bite-01), resulting in a new sentence that describes a “harmless” snake. Not shown in Figure 1 but discussed later, we also introduce an optional faithfulness check as a post-processing step to evaluate the consistency between the transformed graph and the generated sentence, filtering out outputs affected by any eventual parsing or generation errors.
@@
We argue that this combination of symbolic control, neural fluency, and post-hoc verification provides a powerful and flexible approach to sentence meaning transformation, offering a more controllable alternative or complement to LLMs.
@@
To illustrate the practical value of SentenceSmith, we conduct a detailed demonstration study, leveraging our framework to generate hard negative pairs that pose significant challenges for text embedding models. By applying well-defined transformation rules targeting specific semantic phenomena, SentenceSmith creates nuanced sentence pairs that go beyond superficial lexical differences, enabling a more fine-grained evaluation of embedding models. Through this application study, we shed light on their strengths and weaknesses in handling particular linguistic phenomena, addressing a limitation in current benchmarking practices, where such intricacies tend to be obscured.
@@
Broadly speaking, this study contributes to the ongoing effort to pinpoint the limits of Transformer-based Neural Networks’ understanding of sentence meaning. It has been established that even the latest LLMs struggle with important sentence-level meaning phenomena such as negation (parmarLogicBenchSystematicEvaluation2024), and xu2025large finds that LLMs still struggle with exact, syntax-based modes of reasoning, sometimes being outperformed by BART (lewis2019bart), a much smaller and simpler sequence-to-sequence transformer. The fine-grained and dynamic experimental settings enabled by SentenceSmith can shed new light on this central, but still open question regarding the abilities of Transformers.
@@
The remainder of this paper is structured as follows: Section §2 reviews background and related work. Section §3 introduces our neuro-symbolic SentenceSmith framework, detailing its components, including the parser, symbolic graph transformations, generator, and optional faithfulness checker. Section §4 presents an application demonstration, focusing on generating hard negative pairs to challenge state-of-the-art text embedding models and reveal potential weaknesses in their linguistic understanding. Finally, Section §5 concludes the paper, discussing the broader implications of our work and potential directions for future research. We release code and data under a public license.111https://github.com/impresso/sentence-smith
@@ 2 Related Work and Background
The value of using semantic graphs as an intermediate representation, particularly AMR, has been highlighted in two recent surveys (sadeddine-etal-2024-survey; wein-opitz-2024-survey). This approach allows the fusion of neural network power with the expressivity and explicitness of meaning representations, which is especially useful when interpretability and control are required in an application. Related work applies this principle to style transfer (jangra-etal-2022-star), in the MT domain (wein-schneider-2024-lost), and data augmentation (min-etal-2020-syntactic; shou-etal-2022-amr; shou-lin-2023-evaluate; ghosh-etal-2024-abex; kim2024simple). Our work generalizes this approach further and imposes an additional check for validating the faithfulness of the generation; Unlike li-etal-2020-linguistically we do not rely on an inflexible rule-parser and/or LLMs for verification, proving the feasibility of efficiently scaling this approach. Notably, the idea of planning/controlling sentence generation through meaning representation dates decades back (i.a., sondheimer-nebel-1986-logical; mann1986systemic; kasper-1989-flexible; wijnen1990development; bateman1990interfacing), but was limited by inaccuracies in parsing and generation. With stronger parsing and generation systems now available, we argue that effective usage has become feasible.
@@
Text embedding models are crucial for a wide range of NLP tasks, including semantic search, information retrieval, and NLG evaluation (clark2019sentence; muennighoff2022sgpt; gao2023retrieval). Since reimers-gurevych-2019-sentence’s foundational “SBERT” work, multiple branches of embedding model research have emerged. These include enhancing model performance through scaling parameters (wang2023improving) or training data (wang2022texte5), as well as exploring unsupervised embeddings (gao-etal-2021-simcse) and interpretable embeddings (opitz-frank-2022-sbert). Key questions are: What is the accuracy of such embeddings? What level of linguistic understanding resides in those vectors? While the first question is typically addressed through large-scale benchmarks like MTEB (muennighoff-etal-2023-mteb), the second question requires a more fine-grained approach. Moreover, the questions are intertwined, and limits on linguistic understanding might be concealed by averages over large benchmark datasets where individual datasets often have hardly interpretable notions of similarity. We employ SentenceSmith and demonstrate its capacity to generate fine-grained evaluation data that test embedding models’ ability to assess different linguistic phenomena. With special regard to interpretability of text embeddings (opitz2025interpretable), our work is also related to a recent/concurrent line of work by Nastase et al. (nastase-merlo-2024-tracking; nastase-etal-2024-exploring; nastase-etal-2024-exploring-syntactic). In contrast to these works, however, we don’t rely on costly human data creation.
Furthermore, the stasis of most benchmarks has also drawn criticism (fan-etal-2024-nphardeval), and limits evaluation to the available data, with potential ramifications for the trustworthiness of results. By demonstrating how SentenceSmith can generate challenging, trustworthy, and interpretable test sets, we pave the way for more customizable, dynamic and interpretable testing of models.
Paraphrases and building minimal pairs.
How do two texts relate? In a broader context, our work is related to measuring paraphrases and entailment, both long-standing topics of interest in the NLP and ML (bhagat2013paraphrase; bowman-etal-2015-large; zhou-bhat-2021-paraphrase; opitz-etal-2023-amr4nli; gubelmann-etal-2023-truth; NEURIPS2023_575c4500). Instead of building or rating paraphrases, we first and foremost build challenging negatives given a paraphrase set, in a controlled manner, such that the relation between negative and paraphrase is transparent (e.g., negation). This may shed also more light on a problem recently highlighted: Notions of paraphrases empirically differ (michail-etal-2025-paraphrasus) and the differences are not transparent. The controlled manipulations furnished by our SentenceSmith framework allow for more in-depth studies. There are also other efforts that note the value of constructing such relation-controlled minimal linguistic pairs (BLiMP, warstadt-etal-2020-blimp-benchmark; jumelet2025multiblimp)—in contrast to these works, our framework does not require static resources or manual annotation effort.222Concurrent work also used LLMs to similarly generate hard negatives from given paraphrases, with similar application scenario (michail2025examiningmultilingualembeddingmodels; magomere-etal-2025-claims). However, this is a categorically different approach, because no controllable intermediate representation is leveraged, which limits the control over the particular type of relation. Also, the LLM use makes such approaches much more costly, while our approach only depends on smaller neural models.
@@ 3.1 Overview
where are graph transformations that we apply on the graph representation of the input sentence , output of . Finally, generates a text from the final state of the graph.
@@ 3.2 Parameterization
We rely on Abstract Meaning Representation (AMR, banarescu2013abstract) as our graph framework. AMR represents text as directed acyclic graphs, where nodes denote entities and edges capture semantic relations. The overarching goal of AMR is to explicitly encode “Who does what to whom?” This explicitness is a key motivation for our work, as it enables targeted meaning changes within the semantic structure.
@@
As is standard in most AMR applications, we employ off-the-shelf parsing and generation models.333https://github.com/bjascob/amrlib These models, based on pre-trained BART (lewis2019bart), produce linearized sequence graphs from text (parsing) or text from linearized sequence graphs (generation). To facilitate manipulation, we convert the linearized graph into triples of the form <u,r,v>, where u and v are graph nodes, and r is a relation.
@@
The criteria for discarding output sentences depend on the application: If the goal is meaning alteration, we should filter out cases where and mutually entail each other. Conversely, if we seek paraphrases, we should discard cases where contradiction or neutrality is detected. SentenceSmith parameterizes with an efficient NLI-based model from steen-etal-2023-little. This system is robustness-enhanced through data augmentation and achieves strong results on the TRUE faithfulness benchmark (honovich-etal-2022-true-evaluating).444We use https://huggingface.co/juliussteen/DeBERTa-v3-FaithAug.
@@ 4 Application Study
The foil concept is illustrated in Figure 2. The objective is to generate a new sentence that is structurally highly similar to an existing one but semantically distinct. If this transformation is applied to one part of a paraphrase pair, the resulting sentence will not maintain a paraphrase relation with either of the original sentences. In the example, we start with the paraphrases a) A tiger is being bitten by a serpent. and its paraphrastic variant b) The snake bites the tiger. Our parser processes sentence b), and within the generated meaning graph, we perform a semantic role confusion, swapping the agent and patient of the bite event. Finally, our generator produces c) The tiger bites the snake, which is structurally more similar to a) than a) and b) are to each other (i.e., higher lexical overlap). However, the event’s semantics have fundamentally changed, thereby breaking the paraphrase relation.
@@ 4.1.1 Defining Semantic Manipulations
This fundamental manipulation assesses embedding models’ sensitivity to negation. We define negation as altering the polarity of a predicate within an AMR graph. Specifically, this transformation attaches a <n,:polarity,-> edge (in AMR notation, a negation) to a randomly selected node n, excluding pronouns. For example, modifying the predicate approve by adding this edge results in its opposite, disapprove (i.e., in AMR: <n,:polarity,-> is added to node n, where <n,:instance,approve> already exists). The manipulated AMR is then used to generate the negated text. As can be seen in the corresponding example of Figure 3, the truth value of the sentence changes.
RS: Role Swap.
US: Underspecification.
@@
This transformation removes a randomly selected leaf node n from the AMR graph, reducing semantic content, thereby breaking a paraphrase relation. The introduced ambiguity or incompleteness tests how robust embedding models are to partial information.555Note that it can also introduce a polarity change, in case a negation node in the AMR is removed. The notion of “Underspecification” refers to the AMR structure being reduced. In Figure 3, the patient (a snake) has disappeared.
AR: Antonym Replacement
@@
inverts the meaning of a selected node while maintaining the context of the surrounding graph. The system identifies a random non-pronoun node, extracts its lexical stem (e.g., happy in happy-01), and queries WordNet (miller-1993-wordnet) for antonyms. If a suitable antonym is found, the node label is updated while preserving any sense suffix (e.g., -01). This transformation alters meaning more explicitly than polarity negation by directly substituting an opposite term. In the Example for AR given in Figure 3, same as in PN, the truth value changes —while still a biting is happening, this time the agent is small rather than big.
HS: Hypernym Substitution.
@@
This transformation replaces a word with its hypernym, distorting inferences and breaking paraphrase relationships. Consider Penguins can’t fly. If Penguin is replaced with its hypernym Bird, the modified sentence is false (“Birds can’t fly”), demonstrating how hypernym substitution alters meaning while preserving surface structure. Following the same WordNet-based approach as in the antonym replacement, non-pronoun node labels are substituted with randomly selected hypernyms. The node’s graph position and semantic roles remain unchanged, but abstraction increases, disrupting logical inferences. In Figure 3, we observe how snake is replaced with animal. This is also some form of underspecification, but speaking through the AMR, it is achieved by replacing (not removing) a concept (which can also cause a change in truth value, as in the aforementioned bird-example).